OpenStudy (idealist10):

Evaluate the double integral (x+y)dA, where R is the region in the first quadrant bounded by the circles x^2+y^2=4 and x^2+y^2=2y.

OpenStudy (idealist10):

@ParthKohli @radar @Compassionate

ganeshie8 (ganeshie8):

start by sketching the region

OpenStudy (idealist10):

Can we do this problem without using polar coordinates?

ganeshie8 (ganeshie8):

yes buy why do you want to avoid polar ? they're nice..

OpenStudy (idealist10):

I did polar and this is what I got: \[\int\limits_{0}^{\pi/2}\int\limits_{2}^{2\sin \theta}(r \cos \theta+r \sin \theta)r dr d \theta \]

ganeshie8 (ganeshie8):

that looks good, evaluate ?

ganeshie8 (ganeshie8):

work it `dxdy` so that finding the bounds is trivial : ``` y : 0 --> 2 x : sqrt(2y-y^2) --> sqrt(4-y^2) ```

OpenStudy (idealist10):

Thank you!

ganeshie8 (ganeshie8):

yw!